OpenStudy (candy13106):

50+10c=2000+10c solve the equation

OpenStudy (whpalmer4):

Are you sure you copied it correctly?

OpenStudy (whpalmer4):

What happens if you subtract 10c from both sides?

OpenStudy (anonymous):

50+10c=2000+10c I will try to solve your equation. 10c+50=10c+2000(Simplify both sides of the equation) 10c+50+−10c=10c+2000+−10c(Add -10c to both sides) 50=2000 50+−50=2000+−50(Add -50 to both sides) 0=1950

OpenStudy (candy13106):

really

OpenStudy (anonymous):

There are no solutions.

OpenStudy (candy13106):

how

OpenStudy (whpalmer4):

If you made a mistake in copying, there might be a solution. For example, if one of the +10c should be instead -10c there would be a solution

OpenStudy (candy13106):

i copyied it correct

OpenStudy (whpalmer4):

Okay, then there are no solutions.

OpenStudy (candy13106):

ok thanks